OpenStudy (austinl):

Differential Equation, Initial Value Problem I have started to get how to do these fairly easily, however this one has me a tad confused.

OpenStudy (austinl):

\(3y\prime -2y=e^{\frac{-\pi~t}{2}}\) \(y(0)=a\)

OpenStudy (austinl):

I have this so far, \(\Large{y\prime -\frac{2}{3}y=\frac{e^{\frac{\pi t}{2}}}{3}}\)

OpenStudy (loser66):

where is - sign

OpenStudy (austinl):

I forgot it, excuse me.

OpenStudy (loser66):

apply formula, that's it

OpenStudy (austinl):

\(\Large{y\prime -\frac{2}{3}y=\frac{e^{\frac{-\pi t}{2}}}{3}}\) I guess it was just the right side that was confusing me.

OpenStudy (loser66):

rhs is g(t) , just apply formula to get the answer.

OpenStudy (austinl):

\(\Large{y=e^{\frac{2}{3}t}\int e^{-\frac{2}{3}t}~\frac{e^{\frac{\pi t}{2}}}{3}}dt+Ce^{\frac{2}{3}t}\) This look correct?

OpenStudy (loser66):

yep

OpenStudy (loser66):

ok, you are on the right track, just take integral, then replace y (0) to find C , then, plug back

OpenStudy (austinl):

Gimme a minute with this :P

OpenStudy (austinl):

Ok, I have a gnarly answer for the integral.

OpenStudy (austinl):

\(\Large{\frac{2e^{\frac{1}{6}(3\pi -4)t}}{3\pi -4}}\)

OpenStudy (austinl):

Whole answer, \(\Large{y=e^{\frac{2}{3}t}\times\frac{2e^{\frac{1}{6}(3\pi -4)t}}{3\pi -4}+Ce^{\frac{2}{3}t}}\)

OpenStudy (loser66):

look good to me, simplify a little bit

OpenStudy (loser66):

find c , too

OpenStudy (loser66):

@austinL knock knock, finish it, friend. You don't have the final answer yet.

OpenStudy (austinl):

I got this, \(\Large{y(t)=\dfrac{2e^{\frac{1}{6}(3\pi -4)t+\frac{2t}3}}{3\pi-4} +Ce^{\frac{2}{3}t}}\)

OpenStudy (austinl):

\(\Large{y(t)=\dfrac{2e^{\frac{1}{6}(3\pi -4)t+\frac{2t}{3}}}{3\pi -4}+(a-\dfrac{2}{3\pi-4})e^{\frac{2}{3}t}}\) Final answer.

OpenStudy (austinl):

Look ok @amistre64 ?

OpenStudy (amistre64):

\[\Large{y(t)=\dfrac{2e^{\frac{1}{6}(3\pi -4)t+\frac{2t}3}}{3\pi-4} +Ce^{\frac{2}{3}t}}\] \[\Large{y(0)=a=\dfrac{2e^{\frac{1}{6}(0+0)}}{3\pi-4} +Ce^{0}}\] \[\Large{a=\dfrac{2}{3\pi-4} +C}\] as long as the step up to it are fine :) looks ok to me

OpenStudy (austinl):

Very cool, thanks!